The Sir Nigel Gresley Locomotive Trust LtdSubscribe to 60007

Registered Charity Nº1079591


OVERHAUL 2015 - 2019

All the overhaul photos may be found here.



Sir Nigel Gresley is now out of traffic for overhaul. This will probably take three years and cost about £750,000. If you would like to make a donation towards this overhaul please click on one of the the donate buttons.


To donate via PayPal (incl. credit or debit card):Donate         To donate by 'phone: Donate


Notes from information supplied by Darrin Crone, Locomotive Engineer.

Clicking on the images will display a larger version.
Clicking again will close the window. Where an 'X' shows bottom right of the image, click on this (or use the 'f' key) to display full size.
(You may wish to maximise your screen for this)


From April 2018.

23 April 2018

The coach was successfully transported to Cranmore this week, meanwhile losing the York office space, accommodation and storage has involved more work at York. Last week the coach was emptied but this week we have had to sort where we can put everything we need to leave in York. The stuff stored at home has just about been sorted through and some bits have had to go back to York.

On the day of the move the progress of the DB loco was watched with anticipation on Real Time Trains. One of the Engineering Team volunteers had his phone at hand so he kept a watch on the website while a group of us gathered round. About 15 minutes after the booked time the loco was shown to be on its way, much to our relief. After another 15 minutes the DB guard arrived at the North Yard of the NRM.

We had thought that we would have to shunt a WCR 47 in the yard, which was fouling the run round. However, a WCR guard, who fortunately had just popped in to check the 47, and our DB guard discussed the local operating procedures and it was decided that we would propel the coach. This meant we wouldn't need the NRM yard run round, the 47 could stay where it was, and this also meant that we wouldn't have to run round in Holgate sidings.

We left the NRM Yard early and proceeded to Holgate Works without any running round or lengthy stops at signals, so arrived early. The Network Rail guys had set all the point work in their sidings so again we were straight to the traverser, where they had told us to leave the coach when we visited last week.

The next morning our CME, and a member of the Engineering Team volunteers, who had done so much to prepare for the move, and myself went to the Holgate Works to see the loading of the coach. Railway Support Services arrived after a short wait. The coach had already been moved to the road loading point on the traverser and we must thank NR for their cooperation and speedy work.

The ramp was laid out to the end of the trailer and the coach winched on to the low loader by the RSS team without a hitch. CTMS Cranmore were notified that the load was on its way and our SNGLT chairman witnessed its arrival next morning. Work will not start on the coach until summer, so we plan to visit during June and meet with CTMS staff to detail the work that we require.

Earlier in the week the raw cast bronze bogie bearing discs were delivered to the works where they will be machined. At York the existing discs were measured and compared to drawing, just to make sure of dimensions before we start to cut metal on these very expensive parts.

The superheater header delivered last week was stood up and strapped down this week so that the sealing face can be dressed and fitted to the middle cover. Meanwhile the removal of the tender steel work continues around the corridor door at the front of the tender.

Pipework fitting under the cab continues with new pipe being bent up to replace the pipework that ran too close to the Cartazzi wheels. The steam supply pipe to the air pump is also being worked on, annealed, and was cleaned this week.

The ashpan overhaul continues. This week the hopper door had its new cover plate fastened in and the door adjusted to fit. The locking bar was also fitted and the assembly is now very good. The ashpan is to be turned again and the damper screen and door fitted. The diaphragm plate that will locate the ashpan on the loco has been weld repaired where it had been damaged when the old bolting had to be removed by burning.

The insulation for the cylinders was ordered late last week and should have been with us by now. This week we checked round the screws we will need to fix the cladding plates.

The reaming of the horizontal hornsay bolts continued. More than half way round now. The countersink that will be used for these holes has been returned from the tool makers, ground to the correct angle. When not reaming the horizontal hornstay bolts the magdrill is used for the reaming of the loco bogie sidebearer holes. The progress with these has been much slower, however the set up it is much more difficult.

The coupled wheel axleboxes at Grosmont were visited this week by our CME and myself. Measurements were taken on the horizontal journal bearing diameters to confirm the necessity of replacing the horizontal thrust inserts. We also discussed the procedure for their replacement and the machining of the axleboxes with the MPD manager.

The cast bronze discs that fit between the bogie and the loco. They are 26 ½" diameter. 16 April 2018
Photograph: T Crone.

The cast bronze discs

The coach was shunted in the NRM's North Yard so that the DB loco could come straight on to the coach. 17 April 2018
Photograph: Darrin Crone.

In the NRM's North Yard

Not mentioned in the report the reverser reach rod di-pen inspection was completed this week. 18 April 2018
Photograph: Darrin Crone

The reverser reach rod<

In the early evening the DB 66 is seen coupled to the coach ready to leave for the short trip over the network to the Holgate Engineering Works. 18 April 2018
Photograph: Malcolm Bateman/Trevor Camp/SNGLT.

DB 66 coupled to the coach

At the Holgate Engineering Works the coach is winched on to the low loader 19 April 2018.
Photograph: Peter Brackstone.

At Holgate Engineering Works

At Cranmore the coach is winched off the low loader. 20 April 2018
Photograph: Nigel Wilson.

At Cranmore

16 April 2018

The main occupation of the week has been organising the move of the coach from the NRM to the coach overhaulers. As there are so many different parties involved it's been quite time consuming keeping everybody involved informed of any developments. As we stand at the moment the coach is due to move this week. I have to thank the NRM's Chris Graham for his help and support. Chris accompanied me to the Network Rail Holgate Engineering Works after being invited to walk the site by Network Rail. We were shown round and were taken through the procedure to be followed when the coach is brought on to their site, then where it would be later moved to for road loading. The NR personnel are very helpful and didn't give us any feeling of being in their way or it being too much trouble.

To prepare the coach it has now been emptied and its contents gone through and sorted. Much has gone in to store, some in to the NRM workshop, where we will have a further session of sorting through stuff. Quite a bit has been binned and my car was filled with the contents of the office.

The coach dynamo and voltage regulator were loaded in to the guards compartment. They will be refitted by the underframe contractors. The coach gas bottles were also removed.

Meanwhile work continues on the loco frames, tender and boiler. The tender corridor roof from the front of the tender to the rear of the coal space has now been removed. A sketch has been produced based on the LNER drawings and actual measurements of our tender with the intention of getting quotes for new plate.

The bogie frames and wheelsets are receiving varnish.

The first of the loco bogie side bearer holes have been reamed through with the arbour made by an Engineering Team volunteer. It took quite a bit of setting up. It has been found due to the length of the arbour, and the flexibility it introduces, that when going up a size on the reamer that the hole needs to be chamfered with a countersink to stop the reamer from snagging.

As the fireman's side drain cock handle assembly sits on the drop grate bracket that was retrieved from store this week. It was quickly prepared for painting and primed. It was fitted to the loco and the draincock handle put in. Its final alignment will be fixed when the cab is finally fitted after the boiler. New grease nipples were fitted to the draincock cable conduits.

The overhauled whistle and atomiser shut off valve was returned from contractors this week and it's a very good job. That is now in store and is ready to be fitted when the boiler returns. The superheater header and middle coverplate have been returned from machining and these parts will now require final fitting together.

The boiler blanking plates were sent to Llangollen this week, being taken by our official photographer. These will be used when pressure testing the boiler. Our Boiler Project Engineer reports that the copper tubeplate is now fully fixed in place.

The middle crosshead has been di-pen inspected and checked. It has been found to be defect free. The ends of the loco brake shafts have been measured and new bearings can now be manufactured.

Work continues under the cab floor with the pipe fitting team continuing to fit and reconfigure the piping. A hole has been added in the cab floor plate to allow the clear venting of the FVF2 valve which vents vertically downward from its base. Countersunk screws have been fitted in the floor plate where this gives us more space to position pipework.

After making sure of the fit of the coupled hornstays the reaming of the horizontal hornstay bolt holes continued this week. When reaming through the horizontal bolt holes it is very important that this is done when there is no more vertical movement available on the hornstay, as if the hornstay is put back on and the horizontal holes don't line up within a fraction of a thousandth of an inch, the horizontal holes will be damaged by the bolts and the tight fit will be lost.

The middle crosshead is inspected for defects and cracks on 12 April 2018.
Photograph: Richard Swales.

The middle crosshead

The horizontal hornstay bolt holes are reamed. Accuracy and care is needed as the minimum material must be removed. Tony Bickerstaffe operates the magdrill assisted by Roger Turnbull. 12 April 2018
Photograph: Richard Swales.

Horizontal hornstay bolt holes

The superheater header and middle cover returned to York this week after having the sealing faces machined flat. 14 April 2018
Photograph: Darrin Crone.

S'heater header & middle cover

The support coach has been emptied of our office and stores in preparation for its move to contractors for overhaul. 14 April 2018
Photograph: Peter Brackstone.

The support coach

8 April 2018

As weeks 13 and 14 were shortened by the bank holiday weekend, and further shortened by a couple of days break I gave to the volunteers, I decided to report both weeks together.

The superheater header and middle front cover plate have now been skimmed at the works of K D Flavell's and are ready to be returned to York for pressure testing with water, so I'll be looking for volunteers who don't mind getting their feet wet.

Good progress has been made on the bowden cables for the gravity sands and drain cocks. We obtained new stainless steel cables and put them through the conduits. The drain cock ones are now in position. The cab levers and brackets are now painted and the cab drain cock lever assembled. Further bracketry is required before it can be assembled in the cab.

The reverser column and components have been cleaned and prime painted. New ball bearings have been obtained for the assembly. The shaft was assembled back in the column and movement inside of the bearings was measured. We need to know the clearance so we can decide if the bearings need repair or replacement.

The reverser reach rod has been cleaned and initial di-pen examination has revealed no defects. It is in two halves fastened together with fitted bolts. The bolts are still a good tight fit and may be suitable for re-use. There are sharp inside corners at the rod ends and middle joint and these have been carefully cleaned out so that further inspection can be carried out before reassembly.

The assembled clackboxes were returned to York and the handles and squares on the delivery shut off valves were fitted. The squares are fitted with taper pins. It was when operating the handles on the steam valves that it was found that there was some difference in the number of turns required for the valves operation. This was investigated and we found that the steam spindle on one of the valves required some attention. Both valves now operate satisfactorily.

Good progress is being made with the atomiser and whistle valve. New connections are being made, a new valve spindle has been made and a new lever pin fitted. New pipe cones and nuts will be made for the pipes that join the new connections on the valve. The new steam heat valve mounted on the manifold is progressing with the casting of a new valve body.

The tender brakeshaft has now been thoroughly cleaned, finishing around the levers with the wire brush. There is evidence of wastage around the brakeshaft but after inspection it is judged fit for further use. The areas around the levers are yet to be examined for defects.

Further platework from the corridor top has been removed. All the roof has been removed back to the rear of the coal space. There is some more work to be done around the cab door end. Drawings of the tender have been compared to the tender as these drawings will form a basis for ordering the new plate. The cut out plate has now been removed from the workshop and put in the yard with other scrap.

The last piston crosshead has now been thoroughly cleaned and descaled, and is ready for examination. This crosshead looks to be an original 60007 part, whereas the outside ones are from 60026. The outside crossheads have now been examined for defects and measured with a special jig to find the centre of the piston rod taper. From this measurement the slidebar positions can be finalised. All three upper slidebars are now in position on the loco, after the careful alignment against a taught wire through the centre of the cylinder. After calculations are complete new shims will be made to lower or raise the slidebars as necessary to suit the crossheads.

Steady work continues on the ashpan with new plate now put in the bottom around the hopper door opening. The hopper door opening was intact but was thinning at the edges so it was decided to renew it. It is very important to keep the ashpan tight, as if fire can be seen from the outside of the loco it could result in its not being allowed to run on the network. With the plate in place the hole for the door has been re-cut and the door trial fitted.

Work continues on the pipework beneath the cab floor, which stubbornly refuses to go back where it came from. We must thank the pipe fitting team for sticking with it so determinedly.

A number of volunteers went in to the workshop during our break to tidy around our work area and wash the floor. A great job was done and quite a bit of space has been created by tidying up.

We have again visited Llangollen to view progress on the boiler. Along with Llangollen engineering staff, our boiler overhaul engineer showed us around progress so far. The copper tubeplate is now fixed in place all around by patch screws and the fit was closely examined by our CME. It looks to be very good. The re-staying of the back top corners is well advanced. The foundation ring repairs continue and should be finished shortly. We took the opportunity of the visit to deliver more boiler components overhauled at York, including the LNER blow down cock and the gauge frame fittings.

Though the coach still sits at York we are in the process of moving it to its eventual overhaul location. We have had a quote for moving the coach from the north yard to the road loading point in the old carriage works, and a quote for the road move. We have asked a number of traction providers and road hauliers for quotes. The coach requires certain documentation from competent individuals or organisations before it can be moved and this process is ongoing, but approaching completion. The coach has now been examined and brakes tested with a view to getting the certification we need to permit the move.

The bogie frames were moved on to the wheel drop and are now stood next to the bogie wheel sets and axleboxes. The rest of the bogie components have been positioned around the frames ready for the reassembly, however the painting team have stopped the job because they want the frames and wheels varnishing first, and now they want the springs painting for the official works photographs.

The piston rod taper jig in place in the right crosshead. 24 March 201
Photograph: Richard Swales.

The piston rod taper jig

The removal of the corridor plate work from the tender is ongoing. 24 March 201
Photograph: Peter Brackstone.

The corridor plate work

The reverser column has been cleaned tested and prime painted. 24 March 201
Photograph: Peter Brackstone.

The reverser column

A view of the tubeplate from the firebox side. 29 March 2018
Photograph: Richard Swales.

The tubeplate

Our CME closely inspects every inch of the tubeplate joint inside the boiler at Llangollen. 29 March 2018
Photograph: Darrin Crone.

Our CME closely inspects

The bogie frames, wheelsets and suspension components brought together for assembly, if only we could get close enough to put them together but the painting team won't let us anywhere near them during varnishing. 7 April 2018
Photograph: Darrin Crone.

The bogie frames

The middle crosshead is thoroughly cleaned prior to examination for defects and measurement of piston rod taper centre from bearing surfaces. 7 April 2018
Photograph: Darrin Crone.

The middle crosshead
Return to top

From January to March 2016.
From April to June 2016.
From July to September 2016.
From October to December 2016
From January to March 2017.
From April to June 2017.
From July to September 2017.
From October to December 2017.
From January to March 20187.

Page updated on Monday, 23 April 2018   Visitor Counter

Copyright © The Sir Nigel Gresley Locomotive Trust Ltd